# Quantum stochastic thermodynamics in the mesoscopic-leads formulation

### v1

@@ III.2.1 No jump-conditioned evolution
where $B=\Gamma-2F$, $V=iH+\frac{1}{2}B=W-F$. In between jumps, the survival probability $p$—the probability for no jump to occur up to time $t$—evolves under the differential equation
$\begin{split}\frac{\mathrm{d}p}{\mathrm{d}t}&=-p\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\left\langle{L^{\sigma}_{k}}^{\dagger}{L^{\sigma}_{k}}\right\rangle_{r},\\ \end{split}$ (20)
@@ III.2.2 Jump-conditioned updates
Upon recording a jump of type $\sigma_{m}$ in the lead mode $k_{m}$ at time $t_{J_{m}}$, stored in the measurement record as $(t_{J_{m}},k_{m},\sigma_{m})$, $C^{r}$ is updated instantaneously. At $t=t_{J_{m}}$, $\mathrm{d}N_{k,\sigma}(t_{J_{m}})=\delta_{kk_{m}}\delta_{\sigma\sigma_{m}}\in\{0,1\}$ and therefore
$\begin{split}\mathrm{d}C^{r}_{ij}(t_{J_{m}})=&\left\langle G^{+}\left[\rho_{r},{L^{\sigma_{m}}_{k_{m}}}\right]c^{\dagger}_{j}c_{i}\right\rangle_{r}(t_{J_{m}}).\end{split}$ (23)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=\bar{C}^{r}\left(\frac{\mathrm{d}N_{+}(t_{J_{m}})}{\bar{C}^{r}}\right)\bar{C}^{r}.$ (25)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=-C^{r}\left(\frac{\mathrm{d}N_{-}(t_{J_{m}})}{C^{r}}\right)C^{r}.$ (26)
@@
$\begin{split}\left(\frac{\mathrm{d}N_{+}(t)}{\bar{C}^{r}}\right)_{kk}&=\frac{\mathrm{d}N_{k+}(t)}{1-C^{r}_{kk}},\\ \left(\frac{\mathrm{d}N_{-}(t)}{C^{r}}\right)_{kk}&=\frac{\mathrm{d}N_{k-}(t)}{C^{r}_{kk}}.\end{split}$ (27)
@@ III.2.3 Stochastic master equation for the covariance matrix
$\begin{split}\mathrm{d}C^{r}(t)=&\left[-(VC^{r}+C^{r}V^{\dagger})+C^{r}BC^{r}\right]\mathrm{d}t\\ &+\bar{C}^{r}\left(\frac{\mathrm{d}N_{+}(t)}{\bar{C}^{r}}\right)\bar{C}^{r}\\ &-C^{r}\left(\frac{\mathrm{d}N_{-}(t)}{C^{r}}\right)C^{r}.\end{split}$ (28)
@@ IV Bayesian estimation of thermodynamic currents
In the presence of time-dependent driving or multiple reservoirs at different temperatures or chemical potentials, currents of energy and particles will flow through the system, irreversibly producing entropy. The average values of these currents are given by appropriate expectation values with respect to the unconditional density matrix. Importantly, the mesoscopic-leads formalism can reproduce exact NEGF results for these average currents so long as the number of lead modes $L$ is large enough, as shown in previous works Brenes et al. (2020); Lacerda et al. (2023b); Brenes et al. (2023).
In this section, we show how to evaluate the fluctuations of thermodynamic currents along individual trajectories within the mesoscopic-leads formalism. We adopt a Bayesian interpretation of the conditional state, $\rho_{r}$, as the best guess of the system configuration given the set of measurement outcomes. As we will show, the resulting stochastic currents reduce to standard results on average, and yield thermodynamically consistent results for the fluctuations (see Sec. V). To this end, we detail how to express these currents in relation to the trajectory covariance matrix $C^{r}$. In the main text, we present the results assuming perfect detection in all measurement channels. More comprehensive findings considering imperfect detection and different measurement configurations can be found in Appendix C.
It is important to point out that our focus here is on the currents exchanged between the extended system and the residual reservoirs, termed external currents, as opposed to the internal currents exchanged between the system and the leads. We argue that it is the external currents that correspond to those detectable in experiments, especially in the strong coupling regime, where the system hybridises with its environment. For an in-depth discussion of internal and external currents in the mesoscopic-leads formalism, please refer to Refs. Lacerda et al. (2023b, a).
@@ IV.0.1 Particle current
$\begin{split}\frac{\mathrm{d}\langle N\rangle}{\mathrm{d}t}=\sum_{\alpha=1}^{N_{\mathrm{R}}}I_{N_{\alpha}}(t)\,,\end{split}$ (29)
where $I_{N_{\alpha}}(t)=\langle\mathcal{L}_{\alpha}^{+}\left[N\right]\rangle$ denotes the particle current into lead $\alpha$ and $I_{N}(t)=\sum_{\alpha=1}^{N_{\mathrm{R}}}I_{N_{\alpha}}(t)$ is the net particle current. Note that since $\left[H,N\right]=0$, $\mathrm{Tr}\left[N\mathcal{L}_{0}\left[\rho(t)\right]\right]=0$. This can be understood intuitively, as hopping interactions within the extended system conserve the total particle number. Therefore, the total particle current is given soley as the sum over the net particle currents flowing into the extended system from the residual reservoirs.
Similarly, defined through a continuity equation, the net particle current conditioned on the measurement record and thus along a single trajectory, $I^{r}_{N}(t)$, is given by
$\begin{split}\frac{\mathrm{d}\langle N\rangle_{r}}{\mathrm{d}t}=\sum_{\alpha=1}^{N_{\mathrm{R}}}I^{r}_{N_{\alpha}}(t)\,,\end{split}$ (30)
@@
$\begin{split}\left(\Gamma_{\alpha}\right)_{kk},\left(F_{\alpha}\right)_{kk}=\begin{cases}\Gamma_{kk},F_{kk}&\text{if $k$-mode in lead $\alpha$,}\\ 0&\text{otherwise.}\\ \end{cases}\\ \end{split}$ (33)
@@
$\begin{split}\left(\frac{\mathrm{d}N_{+,\alpha}(t)}{\bar{C}^{r}}\right)_{kk}&=\begin{cases}\frac{\mathrm{d}N_{k+,\alpha}(t)}{\bar{C}^{r}_{kk}}&\text{ if $k$-mode in lead $\alpha$,}\\ 0&\text{ otherwise.}\end{cases}\\ \left(\frac{\mathrm{d}N_{-,\alpha}(t)}{C}\right)_{kk}&=\begin{cases}\frac{\mathrm{d}N_{k-,\alpha}(t)}{C^{r}_{kk}}&\text{ if $k$-mode in lead $\alpha$,}\\ 0&\text{ otherwise.}\end{cases}\end{split}$ (34)
The Bayesian nature of the current is apparent in Eq. (32) by the fact that even the extraction of an electron (last term) does not necessarily imply a current of $-1$. This is because the covariance matrix also encompasses our uncertainty about the number of particles within the extended system. Thus, even though a quantum jump definitely corresponds to the detection of an electron, this does not imply that the change in the occupation number of the extended system is $-1$, since initially there was some uncertainty as to the number of electrons in there.
@@
which is consistent with the standard unconditional continuity equation for the particle current, stated in Eq. (29), and is in agreement with results derived in Brenes et al. (2020). The change of the total particle number along a single trajectory in time interval $\left[0,\tau\right]$ conditioned on the measurement record is given by
$\Delta N^{r}(\tau,0)=\int_{0}^{\tau}I^{r}_{N}(t)\mathrm{d}t.$ (37)
@@ IV.0.2 Energy current
$\begin{split}\frac{\mathrm{d}\langle H\rangle}{\mathrm{d}t}&=\sum_{\alpha=1}^{N_{\mathrm{R}}}I_{E_{\alpha}}(t),\\ \end{split}$ (38)
@@
$\begin{split}I^{r}_{E}(t)=&\sum_{\alpha=1}^{N_{\mathrm{R}}}I^{r}_{E_{\alpha}}(t).\\ \end{split}$ (40)
@@ IV.0.3 Heat current and measurement energy
However, in the energy balance along trajectories described by stochastic MEs involving jump operators that are not eigenoperators of the Hamiltonian, a specific term can be singled out Horowitz (2012); Manzano et al. (2015); Levy and Kosloff (2014) that poses ambiguity in its interpretation—whether it should be considered as quantum heat Elouard et al. (2017) or “measurement work” Horowitz (2012); Manzano et al. (2015); Manzano and Zambrini (2022) is not definitively established. That is, Eq. (41) may be split into different contributions
$\Delta E^{r}(\tau,0)=\Delta W^{r}(\tau,0)-\Delta Q^{r}_{\Sigma}(\tau,0)+\Delta E^{r}_{\mathrm{M}}(\tau,0),$ (45)
where $\Delta W^{r}(\tau,0)$ denotes the net work done on the extended system and $\Delta Q^{r}_{\Sigma}$ is the net heat dissipation associated with entropy production in the environment due the exchange of particles and energy between the lead modes and the residual reservoirs, as will be made explicit in the following. We will simply refer to $\Delta E^{r}_{\mathrm{M}}$ as the measurement energy, without further classification. The associated stochastic measurement energy current is given by Manzano and Zambrini (2022)
@@
$\begin{split}\Delta s_{k}&=-\left(\frac{\Delta E_{k}}{T_{r}}-\mu_{r}\frac{\Delta N_{k}}{T_{r}}\right).\end{split}$ (49)
For the conditional measurement energy current for lead $\alpha$ we find
@@
Interestingly, when plugging in the explicit form of the system-lead interaction $H_{\mathrm{int}}=H_{\mathrm{SL}}(t)=\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{L_{\alpha}}\kappa_{\alpha,k}f^{\dagger}_{p_{\alpha}}a_{k}+\mathrm{h.c.}$, as defined in Sec. II, then we find
$E\left[I_{E_{\mathrm{M}}}\right]=-\frac{1}{2}\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{L_{\alpha}}\gamma_{\alpha,k}\langle\kappa_{\alpha,k}f^{\dagger}_{p_{\alpha}}a_{k}+\kappa_{\alpha,k}^{*}a_{k}^{\dagger}f_{p_{\alpha}}\rangle.$ (53)
@@ V Entropy production along a trajectory
In general, the stochastic entropy production along a single trajectory from time 0 to $\tau$ is given by
$S_{\mathrm{tot}}({r_{\left[0,\tau\right]}})=\log\left(\frac{P_{F}(r)}{P_{B}(\tilde{r})}\right),$ (54)
@@ V.1 Two-point measurement scheme
The forward and backward trajectories can be precisely defined within the framework of the “two-point measurement” (TPM) scheme, as schematically shown in Fig. 3 and detailed in Ref. Manzano and Zambrini (2022). In the TPM scheme, the (stochastic) time-evolution of the state in the open interval $(0,\tau)$ is framed by two measurements, one at the start ($t=0$) and one at the end of each trajectory ($t=\tau$). The full trajectory in the closed interval $\left[0,\tau\right]$, including the projective measurements, is termed the forward trajectory. The observables measured here in the projective measurements at the beginning and end of each trajectory (see Fig. 3) are time-reversal invariant. In particular, we consider the case in which each of the projective measurements in the time-reversed backward process produces the same outcome as in the forward process. This means that the measurement record of the backward trajectory in between the two measurements $\tilde{r}_{\left(0,\tau\right)}$ is exactly the time-reversed measurement record of the forward trajectory $r_{\left(0,\tau\right)}$ with $N_{j}$ recorded jumps, so $\tilde{r}_{\left(0,\tau\right)}=\{(\tau-t_{J_{N_{j}}},k_{N_{j}},-\sigma_{N_{j}}),(\tau-t_{J_{N_{j}-1}},k_{N_{j}-1},-\sigma_{N_{j}-1}),\dots,(\tau-t_{J_{1}},k_{1},-\sigma_{1})\}$ Manzano et al. (2018).
@@
$S_{\mathrm{tot}}({r_{\left[0,\tau\right]}})=\log\left(\frac{p^{0}_{n}}{p^{\tau}_{m}}\right)-\Sigma_{r_{\left(0,\tau\right)}}.$ (57)
@@
where $\Delta E_{k,\alpha}$ is assumed to be the energy of the lead mode $k$ in lead $\alpha$, as discussed also Sec. IV.0.3. A discussion of entropy production and fluctuation theorems accounting for imperfect detection may be found in Ref. Surace and Tagliacozzo (2022).
@@ V.2 Efficient computation of $p^{0}_{n}$, $C^{r}(0)$ and $p^{\tau}_{m}$
|$\ket{s_{n}}=U_{s}\ket{0},\hskip 28.45274ptU_{s}=\prod_{i=1}^{N_{\mathrm{S}}+L}\left(d_{i}^{\dagger}d_{i}\right)^{s_{n}^{i}},$ (62)
@@
and therefore $U_{s}$ is Gaussian.
The probability of projecting $\rho(0)$ onto its eigenstate |$\ket{s_{n}}$ during the first projective measurement is determined by the corresponding eigenvalue $p^{0}_{n}$.
@@ V.2.2 Covariance matrix after the first projective measurement
where $\Lambda=\mathrm{diag}(\mathrm{eigenvals(C(0))})$.
The covariance matrix $C^{r}(0)$, computed with respect to the eigenstate |$\ket{s_{n}}$ of $\rho(0)$, is given by
@@
where $(n_{s_{n}})_{ij}=\bra{s_{n}}d_{j}^{\dagger}d_{i}\ket{s_{n}}$, as detailed in Appendix A. Note that the set of matrices $\{n_{s_{i}}\}$ is simply given by the set of all $2^{N_{\mathrm{S}}+L}$ diagonal matrices with bit-strings of length $N_{\mathrm{S}}+L$ (with letters either 0 or 1) on their diagonal. For instance, for $N_{\mathrm{S}}+L=2$: $\{n_{s_{i}}\}=\{\mathrm{diag}\left((0,0)\right),\mathrm{diag}\left((0,1)\right),\mathrm{diag}\left((1,0)\right),\mathrm{diag}\left((1,1)\right)\}$.

### v2

@@ III.2.1 No jump-conditioned evolution
where $B=\Gamma-2F$, $V=iH+\frac{1}{2}B=W-F$. In between jumps, the survival probability pp—the probability for no jump to occur up to time tt—evolves under the differential equation
[…]
@@ III.2.2 Jump-conditioned updates
Upon recording a jump of type $\sigma_{m}$ in the lead mode $k_{m}$ at time $t_{J_{m}}$, stored in the measurement record as $(t_{J_{m}},k_{m},\sigma_{m})$, $C^{r}$ is updated instantaneously. At $t=t_{J_{m}}$, $\mathrm{d}N^{\sigma}_{k}(t_{J_{m}})=\delta_{kk_{m}}\delta_{\sigma\sigma_{m}}\in\{0,1\}$ and therefore
$\begin{split}\mathrm{d}C^{r}_{ij}(t_{J_{m}})=&\left\langle\mathcal{G}^{+}\left[\rho_{r},{L^{\sigma_{m}}_{k_{m}}}\right]c^{\dagger}_{j}c_{i}\right\rangle_{r}(t_{J_{m}}).\end{split}$ (23)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=\bar{C}^{r}\left(\frac{\mathrm{d}N^{+}(t_{J_{m}})}{\bar{C}^{r}}\right)\bar{C}^{r}.$ (25)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=-C^{r}\left(\frac{\mathrm{d}N^{-}(t_{J_{m}})}{C^{r}}\right)C^{r}.$ (26)
@@
$\begin{split}\left(\frac{\mathrm{d}N^{+}(t)}{\bar{C}^{r}}\right)_{kk}&=\frac{\mathrm{d}N^{+}{k}(t)}{1-C^{r}_{kk}},\\ \left(\frac{\mathrm{d}N^{-}(t)}{C^{r}}\right)_{kk}&=\frac{\mathrm{d}N^{-}_{k}(t)}{C^{r}_{kk}}.\end{split}$ (27)
@@ III.2.3 Stochastic master equation for the covariance matrix
$\begin{split}\mathrm{d}C^{r}(t)=&\left[-(VC^{r}+C^{r}V^{\dagger})+C^{r}BC^{r}\right]\mathrm{d}t\\ &+\bar{C}^{r}\left(\frac{\mathrm{d}N^{+}(t)}{\bar{C}^{r}}\right)\bar{C}^{r}\\ &-C^{r}\left(\frac{\mathrm{d}N^{-}(t)}{C^{r}}\right)C^{r}.\end{split}$ (28)
@@ IV Bayesian estimation of thermodynamic currents
In the presence of time-dependent driving or multiple reservoirs at different temperatures or chemical potentials, currents of energy and particles will flow through the system, irreversibly producing entropy. The average values of these currents are given by appropriate expectation values with respect to the unconditional density matrix. Importantly, the mesoscopic-leads formalism can reproduce exact NEGF results for these average currents so long as the number of lead modes $L$ is large enough, as shown in previous works Brenes et al. (2020); Lacerda et al. (2023a); Brenes et al. (2023). Note that the currents of interest in our framework are those exchanged between the lead modes and the residual reservoir. This distinction is important, as we argue that the currents measured in mesoscopic experiments, e.g. by an ammeter, correspond more closely to these internal reservoir currents—especially in the strong-coupling regime. In this regime, the presence of coherences and hybridization blurs the boundary between the system and its environment, making a clear separation between them effectively impossible. In this work, we aim to capture the fluctuations of the time-averaged current that would be measured in an experimental setting. These experimentally accessible fluctuations can be obtained from our theoretical framework by analyzing the statistics of the integrated charge along quantum trajectories.
In this section, we evaluate fluctuations of thermodynamic currents along individual trajectories using the mesoscopic-leads formalism. We interpret the conditional state $\rho_{r}$ as the best estimate of the system’s state, given the measurement record. The current is computed with respect to this conditional state, which evolves as new data is acquired, reflecting a Bayesian update of the observer’s knowledge, as commonly understood in the theory of quantum trajectories Wiseman and Milburn (2009). As we will show, the resulting stochastic currents reduce to standard results on average, and yield thermodynamically consistent results for the fluctuations (see Sec. V). To this end, we detail how to express these currents in relation to the trajectory covariance matrix $C^{r}$. In the main text, we present the results assuming perfect detection in all measurement channels. More comprehensive findings considering imperfect detection and different measurement configurations can be found in Appendix C.
It is important to point out that our focus here is on the currents exchanged between the extended system and the residual reservoirs, termed external currents, as opposed to the internal currents exchanged between the system and the leads. We argue that it is the external currents that correspond to those detectable in experiments, especially in the strong coupling regime, where the system hybridises with its environment. For an in-depth discussion of internal and external currents in the mesoscopic-leads formalism, please refer to Refs. Lacerda et al. (2023a, 2024).
@@ IV.0.1 Particle current
[…]
where $I_{N_{\alpha}}(t)=\langle\mathcal{L}_{\alpha}^{+}\left[N\right]\rangle$ denotes the particle current into lead $\alpha$ and $I_{N}(t)=\sum_{\alpha=1}^{N_{\mathrm{R}}}I_{N_{\alpha}}(t)$ is the net particle current. Note that since $\left[H,N\right]=0$, $\mathrm{Tr}\left[N\mathcal{L}_{0}\left[\rho(t)\right]\right]=0$. This can be understood intuitively, as hopping interactions within the extended system conserve the total particle number. Therefore, the total particle current is given soley as the sum over the net particle currents flowing into the extended system from the residual reservoirs. Similarly, defined through a continuity equation, the net particle current conditioned on the measurement record and thus along a single trajectory, $I^{r}_{N}(t)$, is given by
[…]
@@
[…]
@@
[…]
The Bayesian nature of the current is apparent in Eq. (32) by the fact that even the extraction of an electron (last term) does not necessarily imply a current of 1-1. This is because the covariance matrix also encompasses our uncertainty about the number of particles within the extended system prior to a detection event. Thus, even though a quantum jump definitely corresponds to the detection of an electron, this does not imply that the change in the occupation number of the extended system is 1-1, since initially there was some uncertainty as to the number of electrons in there.
@@
which is consistent with the standard unconditional continuity equation for the particle current, stated in Eq. (29), and is in agreement with results derived in Brenes et al. (2020).
@@ IV.0.2 Energy current
$\begin{split}\frac{\mathrm{d}\langle H\rangle}{\mathrm{d}t}&=\sum_{\alpha=1}^{N_{\mathrm{R}}}I_{E_{\alpha}}(t),\end{split}$ (37)
@@
$\begin{split}I^{r}_{E}(t)=&\sum_{\alpha=1}^{N_{\mathrm{R}}}I^{r}_{E_{\alpha}}(t).\end{split}$ (39)
@@ IV.0.3 Heat current and measurement energy
However, in the energy current balance along trajectories described by stochastic MEs involving jump operators that are not eigenoperators of the Hamiltonian, a specific term can be singled out Horowitz (2012); Manzano et al. (2015); Levy and Kosloff (2014) that poses ambiguity in its interpretation—whether it should be considered as “quantum heat” current Elouard et al. (2017) or “measurement work” currentHorowitz (2012); Manzano et al. (2015); Manzano and Zambrini (2022) is not definitively established. It is given by Manzano and Zambrini (2022)
@@
$\begin{split}\Delta s_{k}&=-\left(\frac{\Delta E_{k}}{T_{r}}-\mu_{r}\frac{\Delta N_{k}}{T_{r}}\right),\end{split}$ (47)
where $T_{r}$ and $\mu_{r}$ are the temperature and chemical potential, respectively, of the residual reservoir that mode $k$ interacts with. For the conditional measurement energy current for lead $\alpha$ we find
@@
Interestingly, when plugging in the explicit form of the system-lead interaction $H_{\mathrm{int}}=H_{\mathrm{SL}}(t)=\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{l_{\alpha}}\kappa_{\alpha,k}d^{\dagger}_{p_{\alpha}}a_{k}+\mathrm{h.c.}$, as defined in Sec. II, then we find
$E\left[I_{E_{\mathrm{M}}}\right]=-\frac{1}{2}\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{l_{\alpha}}\gamma_{\alpha,k}\langle\kappa_{\alpha,k}d^{\dagger}_{p_{\alpha}}a_{k}+\kappa_{\alpha,k}^{*}a_{k}^{\dagger}f_{p_{\alpha}}\rangle.$ (51)
@@ V Entropy production along a trajectory
In general, the stochastic entropy production in the extended system along a single trajectory from time 0 to $\tau$ is given by
[…]
@@ V.1 Two-point measurement scheme
The forward and backward trajectories can be precisely defined within the framework of the “two-point measurement” (TPM) scheme, as schematically shown in Fig. 3 and detailed in Ref. Manzano and Zambrini (2022). In the TPM scheme, the (stochastic) time-evolution of the state in the open interval $(0,\tau)$, yielding the measurement record ${r}_{\left(0,\tau\right)}$ is framed by two measurements, one at the start ($t=0$) and one at the end of each trajectory ($t=\tau$). The full trajectory in the closed interval $\left[0,\tau\right]$, including the projective measurements, is termed the forward trajectory and we denote the corresponding measurement record by ${r}_{\left[0,\tau\right]}$. The observables measured here in the projective measurements at the beginning and end of each trajectory (see Fig. 3) are time-reversal invariant. In particular, we consider the case in which each of the projective measurements in the time-reversed backward process produces the same outcome as in the forward process. This means that the measurement record of the backward trajectory in between the two measurements $\tilde{r}_{\left(0,\tau\right)}$ is exactly the time-reversed measurement record of the forward trajectory $r_{\left(0,\tau\right)}$ with $N_{j}$ recorded jumps, so $\tilde{r}_{\left(0,\tau\right)}=\{(\tau-t_{J_{N_{j}}},k_{N_{j}},-\sigma_{N_{j}}),(\tau-t_{J_{N_{j}-1}},k_{N_{j}-1},-\sigma_{N_{j}-1}),\dots,(\tau-t_{J_{1}},k_{1},-\sigma_{1})\}$ Manzano et al. (2018).
@@
$S_{\mathrm{tot}}({r_{\left[0,\tau\right]}})=\log\left(\frac{p^{0}_{n}}{p^{\tau}_{m}}\right)+\Sigma_{r_{\left(0,\tau\right)}}.$ (55)
@@
where $\Delta E_{k,\alpha}$ is assumed to be the energy of the lead mode $k$ in lead $\alpha$, as discussed also Sec. IV.0.3. A discussion of entropy production and fluctuation theorems accounting for imperfect detection may be found in Ref. Ferri-Cortés et al. (2023).
@@ V.2 Efficient computation of $p^{0}_{n}$, $C^{r}(0)$ and $p^{\tau}_{m}$
[…]
@@
[…]
@@ V.2.2 Covariance matrix after the first projective measurement
[…]
@@
[…]